OpenStudy (anonymous):

divide. write the quotients in lowest terms. 13x/5y divided by 2x.

OpenStudy (anonymous):

(13x/5y)/2x = (13x/5y) * (1/2x) = 13x/10xy = 13/10y

OpenStudy (diyadiya):

oops yeah :D

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

13x/5y = (2x)(6/5y)+x/5y so quotient = 6/5y